TIFFCP(1) TIFFCP(1) NAME tiffcp−copy(andpossiblyconvert)aTIFFfile SYNOPSIS tiffcp[options]src1.tif...srcN.tifdst.tif DESCRIPTION tiffcpcombinesoneormorefilescreatedaccordingtotheTagImageFileFormat,Revision6.0 intoasingleTIFFfile.Becausetheoutputfilemaybecompressedusingadifferentalgorithm thantheinputfiles,tiffcpismostoftenusedtoconvertbetweendifferentcompressionschemes. Bydefault,tiffcpwillcopyalltheunderstoodtagsinaTIFFdirectoryofaninputfiletotheassociateddirectoryintheoutputfile. tiffcpcanbeusedtoreorganizethestoragecharacteristicsofdatainafile,butitisexplicitly intendedtonotalterorconverttheimagedatacontentinanyway. OPTIONS −BForceoutputtobewrittenwithBig-Endianbyteorder.Thisoptiononlyhasaneffect whentheoutputfileiscreatedoroverwrittenandnotwhenitisappendedto. −CSuppresstheuseof‘‘stripchopping’’whenreadingimagesthathaveasinglestrip/tileof uncompresseddata. −cSpecifythecompressiontousefordatawrittentotheoutputfile:nonefornocompression,packbitsforPackBitscompression,lzwforLempel-Ziv&Welchcompression,jpeg forbaselineJPEGcompression,zipforDeflatecompression,g3forCCITTGroup3 (T.4)compression,andg4forCCITTGroup4(T.6)compression.Bydefaulttiffcpwill compressdataaccordingtothevalueoftheCompressiontagfoundinthesourcefile. TheCCITTGroup3andGroup4compressionalgorithmscanonlybeusedwithbilevel data. Group3compressioncanbespecifiedtogetherwithseveralT.4-specificoptions:1dfor 1-dimensionalencoding,2dfor2-dimensionalencoding,andfilltoforceeachencoded scanlinetobezero-filledsothattheterminatingEOLcodeliesonabyteboundary. Group3-specificoptionsarespecifiedbyappendinga‘‘:’’-separatedlisttothe‘‘g3’’option; e.g.−cg3:2d:filltoget2D-encodeddatawithbyte-alignedEOLcodes. LZWcompressioncanbespecifiedtogetherwithapredictorvalue.Apredictorvalueof2 causeseachscanlineoftheoutputimagetoundergohorizontaldifferencingbeforeitis encoded;avalueof1forceseachscanlinetobeencodedwithoutdifferencing.LZW-spe-cificoptionsarespecifiedbyappendinga‘‘:’’-separatedlisttothe‘‘lzw’’option;e.g.−c lzw:2forLZWcompressionwithhorizontaldifferencing. −fSpecifythebitfillordertouseinwritingoutputdata.Bydefault,tiffcpwillcreateanew filewiththesamefillorderastheoriginal.Specifying−flsb2msbwillforcedatatobe writtenwiththeFillOrdertagsettoLSB2MSB,while−fmsb2lsbwillforcedatatobe writtenwiththeFillOrdertagsettoMSB2LSB. −lSpecifythelengthofatile(inpixels).tiffcpattemptstosetthetiledimensionssothat nomorethan8kilobytesofdataappearinatile. −LForceoutputtobewrittenwithLittle-Endianbyteorder.Thisoptiononlyhasaneffect whentheoutputfileiscreatedoroverwrittenandnotwhenitisappendedto. −MSuppresstheuseofmemory-mappedfileswhenreadingimages. −pSpecifytheplanarconfigurationtouseinwritingimagedatathathasone8-bitsample perpixel.Bydefault,tiffcpwillcreateanewfilewiththesameplanarconfigurationas theoriginal.Specifying−pcontigwillforcedatatobewrittenwithmulti-sampledata packedtogether,while−pseparatewillforcesamplestobewritteninseparateplanes. January9,1996 1 TIFFCP(1) TIFFCP(1) −rSpecifythenumberofrows(scanlines)ineachstripofdatawrittentotheoutputfile.By default,tiffcpattemptstosettherows/stripthatnomorethan8kilobytesofdataappear inastrip. −sForcetheoutputfiletobewrittenwithdataorganizedinstrips(ratherthantiles). −tForcetheoutputfiletobewrittenwtihdataorganizedintiles(ratherthanstrips). optionscanbeusedtoforcetheresultantimagetobewrittenasstripsortilesofdata, respectively. −wSpecifythewidthofatile(inpixels).tiffcpattemptstosetthetiledimensionssothatno morethan8kilobytesofdataappearinatile. EXAMPLES ThefollowingconcatenatestwofilesandwritestheresultusingLZWencoding: tiffcp-clzwa.tifb.tifresult.tif ToconvertaG31d-encodedTIFFtoasinglestripofG4-encodeddatathefollowingmightbe used: tiffcp-cg4-r10000g3.tifg4.tif (1000isjustanumberthatislargerthanthenumberofrowsinthesourcefile.) SEEALSO pal2rgb(1),tiffinfo(1),tiffcmp(1),tiffmedian(1),tiffsplit(1),libtiff(3) January9,1996 2